McCullough Hospital uses a job-order costing system to assign costs to its patients. Its direct materials include a variety of items such as pharmaceutical drugs, heart valves, artificial hips, and pacemakers. Its direct labor costs (e.g., surgeons, anesthesiologists, radiologists, and nurses) associated with specific surgical procedures and tests are traced to individual patients. All other costs, such as depreciation of medical equipment, insurance, utilities, incidental medical supplies, and the labor costs associated with around-the-clock monitoring of patients are treated as overhead costs.

Historically, McCullough has used one predetermined overhead rate based on the number of patient-days (each night that a patient spends in the hospital counts as one patient-day) to allocate overhead costs to patients. Recently a member of the hospital’s accounting staff has suggested using two predetermined overhead rates (allocated based on the number of patient-days) to improve the accuracy of the costs allocated to patients. The first overhead rate would include all overhead costs within the Intensive Care Unit (ICU) and the second overhead rate would include all Other overhead costs. Information pertaining to the hospital’s estimated number of patient-days, its estimated overhead costs, and two of its patients—Patient A and Patient B—is provided below:

ICU Other Total
Estimated number of patient-days 3,600 32,400 36,000
Estimated fixed overhead cost $ 5,954,400 $ 32,756,400 $ 38,710,800
Estimated variable overhead cost per patient-day $ 290 $ 123
Patient A Patient B
Direct materials $ 6,100 $ 7,800
Direct labor $ 29,000 $ 39,200
Total number of patient-days (including ICU) 25 30
Number of patient-days spent in ICU 0 18

Required:

1. Assuming McCullough calculates two overhead rates as recommended by the staff accountant, calculate:

  1. The ICU and Other overhead rates.
  2. The total cost, including direct materials, direct labor and applied overhead, assigned to Patient A and Patient B.

2. Assuming McCullough continues to use only one predetermined overhead rate, calculate:

  1. The predetermined overhead rate.
  2. The total cost, including direct materials, direct labor and applied overhead, assigned to Patient A and Patient B.

(Round "Predetermined overhead rate" to 2 decimal places. Round other intermediate calculations and final answers to the nearest dollar amount.)

1a. Predetermined ICU overhead rate per patient-day
Predetermined Other overhead rate per patient-day
1b. Total cost for patient A
Total cost for patient B
2a. Predetermined overhead rate per patient-day
2b. Total cost for patient A
Total cost for patient B
Answer #1

1a) Predetermine ICU overhead rate = (5954400/3600+290) = 1944 per patient days

Predetermine other overhead rate = (32756400/32400+123) = 1134 per patient days

1b) Total Cost

Patient A Patient B
Direct material 6100 7800
Direct labor 29000 39200
ICU overhead 0 18*1944 = 34992
Other overhead 25*1134 = 28350 12*1134 = 13608
Total 63450 95600

2a) Predetermine overhead rate = (3600*290+32400*123+38710800)/36000 = 1215

2b) Total Cost

Patient A Patient B
Direct material 6100 7800
Direct labor 29000 39200
overhead 25*1215 = 30375 30*1215 = 36450
Total 65475 83450